Sequential Method

Jasmine Company manufactures both pesticide and liquid fertilizer, with each product manufactured in separate departments. Three support departments support the production departments: Power, General Factory, and Purchasing. Budgeted data on the five departments are as follows:


Support Departments
Producing Departments
Power General
Factory
Purchasing Pesticide Liquid
Fertilizer
Overhead $80,000 $312,000    $169,000    $78,800 $107,700
Square feet 1,500 —    1,500    4,200 4,800
Machine hours 1,403    1,345    24,000 8,000
Purchase orders 20    40    —    120 60

The company does not break overhead into fixed and variable components. The bases for allocation are power—machine hours; general factory—square feet; and purchasing—purchase orders.

The company has decided to use the sequential method of allocation instead of the direct method. The support departments are ranked in order of highest cost to lowest cost.

Required:

1. Allocate the overhead costs to the producing departments using the sequential method. Carry out allocation ratios to four decimal places. Use these numbers for subsequent calculations. Round allocated costs to the nearest dollar. If an amount is zero, enter "0".

Allocation ratios:

Power General Factory Purchasing Pesticide Liquid Fertilizer
Square feet
Machine hours
Purchase orders

Cost allocation:

Power General Factory Purchasing Pesticide Liquid Fertilizer
Direct costs $ $ $ $ $
General Factory               
Purchasing            
Power         
Total $ $ $ $ $

2. Using machine hours, compute departmental overhead rates. (Round the overhead rates to the nearest cent.)

Overhead Rates
Pesticide $ per machine hour
Liquid Fertilizer $ per machine hour
Answer #1

Answer

  • [1]

Power

General Factory

Purchasing

Pesticide

Liquid Fertilizer

Square feet [1500 + 1500 + 4200 + 4800 = 12000]

0.1250

0.1250

0.3500

0.4000

Machine hours [24000 + 8000 = 32000]

0.7500

0.2500

Purchase orders [ 20 + 120 + 60 = 200]

0.1000

0.6000

0.3000

--Working

Power

General Factory

Purchasing

Pesticide

Liquid Fertilizer

Square feet [1500 + 1500 + 4200 + 4800 = 12000]

=1500/12000

=1500/12000

=4200/12000

=4800/12000

Machine hours [24000 + 8000 = 32000]

=24000/32000

=8000/32000

Purchase orders [ 20 + 120 + 60 = 200]

=20/200

=120/200

=60/200

  • [2]

Power

General Factory

Purchasing

Pesticide

Liquid Fertilizer

Direct costs

$80,000

$312,000

$169,000

$78,800

$107,700

General Factory

$39,000

($312,000)

$39,000

$109,200

$124,800

Purchasing

$20,800

($208,000)

$124,800

$62,400

Power

($139,800)

$104,850

$34,950

Total

$0

$0

$0

$417,650

$329,850

--Working

Power

General Factory

Purchasing

Pesticide

Liquid Fertilizer

Direct costs

80000

312000

169000

78800

107700

General Factory

=312000*0.125

-312000

=312000*0.125

=312000*0.35

=312000*0.4

Purchasing

=208000*0.1

=-169000-39000

=208000*0.6

=208000*0.3

Power

=-80000-39000-20800

=139800*0.75

=139800*0.25

  • [3]

Overhead Rates

Pesticide

$                       17.40

$ per machine hour

Liquid Fertilizer

$                       41.23

$ per machine hour